ROTATIONAL SPECTRA AND MOLECULAR CONFORMATION OF TWO CONFORMERS OF LEUCI-NAMIDE

\underline{ANDREW R. CONRAD} and MICHAEL J. TUBERGEN, Department of Chemistry, Kent State University, Kent, Ohio 44242; RICHARD J. LAVRICH, Department of Chemistry and Biochemistry, College of Charleston, 66 George Street, Charleston, South Carolina 29424.

The rotational spectra of the parent species of two conformers of the amino acid derivative leucinamide have been recorded between 10.5 and 20 GHz using a Balle-Flygare Fourier-transform microwave spectrometer. 6 a- and b-type rotational transitions were recorded for conformer I and 23 a- and b-type rotational transitions were recorded for conformer II. The rotational constants are A = 2274.94(9) MHz, B = 1033.336(7) MHz, and C = 911.732(3) MHz and A = 2752.775(8) MHz, B = 843.502(1) MHz, and C = 796.721(1) MHz for conformers I and II respectively. The congested hyperfine patterns arising from the two ^{14}N nuclei were not resolved. Comparison of the experimental moments of inertia to those derived from the lowest energy ab initio (MP2/6-311++G(d,p)) structures indicates gas-phase backbone structures that differ from both the crystal structures of leucinamide^{a} and the gas-phase structures of the amino acid leucine^{b}.

aP. Belton and Y. Wang, J. Mol. Struct., 602-603, 2002, 71-78.

bE. J. Cocinero, A. Lesarri, J.-U. Grabow, J. C. López, and J. L. Alonso, ChemPhysChem, 8, 2007, 599-604.